1.    Eight kilometers below the surface of the ocean the pressure is 82Mpa. Determine the density of the seawater at this depth if the density at the surface is 1025kg/m3 and the average bulk modulus of elasticity is 2.3GPa.



1
Expert's answer
2021-08-16T17:31:21-0400

Gives

h=8km

Bulk modulus

"\beta=\frac{∆pV}{∆V}"

"\frac{∆p}{\beta}=\frac{∆V}{V}"

We know that

"\frac{∆V}{V}=\frac{\frac{m}{\rho}-\frac{m}{\rho'}}{\frac {m}{\rho}}"

"\frac{∆V}{V}={1}-{\frac{\rho}{\rho'}}"

"\frac{∆p}{\beta}=1-\frac{\rho}{\rho'}"

"\rho'=\frac{\rho}{1-\frac{∆p}{\beta}}"

Put value

"\rho'=\frac{1025}{1-\frac{80.36\times10^6-82\times10^6}{2.3\times10^9}}"

"\rho'=\frac{1025}{1-\frac{1.64\times10^6}{2.3\times10^9}}"

"\rho'=\frac{1025}{1-7.13\times10^{-4}}"

"\rho'=1025.73kg/m^3"